# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v1

@@ Abstract
In real-world scenarios, scanned point clouds are often incomplete due to occlusion issues. The task of self-supervised point cloud completion involves reconstructing missing regions of these incomplete objects without the supervision of complete ground truth. Current self-supervised methods either rely on multiple views of partial observations for supervision or overlook the intrinsic geometric similarity that can be identified and utilized from the given partial point clouds. In this paper, we propose MAL-SPC, a framework that effectively leverages both object-level and category-specific geometric similarities to complete missing structures. Our MAL-SPC does not require any 3D complete supervision and only necessitates a single partial point cloud for each object. Specifically, we first introduce a Pattern Retrieval Network to retrieve similar position and curvature patterns between the partial input and the predicted shape, then leverage these similarities to densify and refine the reconstructed results. Additionally, we render the reconstructed complete shape into multi-view depth maps and design an adversarial learning module to learn the geometry of the target shape from category-specific single-view depth images. To achieve anisotropic rendering, we design a density-aware radius estimation algorithm to improve the quality of the rendered images. Our MAL-SPC yields the best results compared to current state-of-the-art methods. We will make the source code publicly available at https://github.com/ltwu6/malspc.
@@ I Introduction
Point clouds, which represent 3D instances as a collection of points in 3D space, have attracted increasing interest due to the growing popularity of 3D scanning devices such as LiDAR and depth cameras. Point clouds offer a detailed representation of the geometric features and spatial structures of real-world 3D objects and scenes. They are widely applied in various fields, such as autonomous driving[1, 2, 3, 4, 5], motion analysis[6, 7], and augmented reality[8, 9]. Accurate and complete shape representation plays a vital role in understanding 3D point clouds. However, point clouds scanned in practical scenarios are often incomplete due to occlusion. This incompleteness severely impacts the performance of downstream tasks, such as 3D object detection [10, 11, 12] and scene understanding [13, 14].
@@
To solve this problem, point cloud completion methods have emerged. Point cloud completion aims to infer the missing parts of an object based on the incomplete point cloud. Most of the existing point cloud completion methods rely on a large number of complete point clouds as supervision [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, in real-world scenes, clean and complete point clouds are really few since current devices struggle to directly capture such data.
@@
To overcome these limitations, some research has focused on unsupervised learning approaches for point cloud completion. Among these methods, some employ complete shapes unpaired with the partial input as priors to drive the completion process [25, 26, 27, 28, 29, 30, 31], while others utilize 2D signals such as depth maps or silhouette images from multiple viewpoints as supervision [32, 33]. However, the former still requires complete point clouds to acquire shape priors, and the latter relies on 2D images of objects and the corresponding projection parameters. Recently, self-supervised methods, which only require partial inputs for supervision, have been proposed. These methods estimate the missing parts of the point cloud by utilizing the multi-view consistency of partial scans and the robust learning capabilities of the masked models. However, they face some limitations. First, these methods rely on multi-view partial point clouds of each object as shape constraints or as training data to drive the shape learning process. Nevertheless, obtaining point clouds from multiple viewpoints is also impractical in real-world applications. Secondly, they fail to utilize the inherent local similarities and category-specific geometric properties of objects.
@@
From our perspective, it is commonly observed that an object frequently exhibits similar or repetitive local structures. For instance, a table typically features four identical cylindrical legs. This observation implies that the existing and missing sections of a point cloud are likely to share similar local geometric patterns. Moreover, objects within the same category generally display consistent structures. For example, most airplanes feature two flat wings. Consequently, these inherent geometric similarities, specific to each category, can be utilized to guide the completion of missing regions in point clouds.
@@
In this paper, we propose MAL-SPC, a Multi-view Adversarial Learning framework for Self-supervised Point cloud Completion. One of the core components, the Pattern Retrieval Network (PRN), reconstructs complete shapes by initially generating a coarse shape and subsequently encoding local geometric patterns from both the coarse and partial input shapes. These patterns manifest in two forms: position encodings, quantified as the average distance between each point and its nearest neighbors, and curvature encodings, defined by the variations in normals between each point and its closest neighbors. Subsequently, a cross-attention-based module is introduced to retrieve correlated local patterns for the coarse shape from the partial shape. These retrieved patterns are further utilized to guide the densification and refinement of the coarse shape, capitalizing on the object-level geometric similarities and repetitiveness inherent in the data.
@@
In addition, we introduce a Multi-view Adversarial Network (MAN) to learn the shared geometry of each category. Objects within the same category commonly exhibit remarkably similar shape distributions because they share many typical geometric structures. This similarity suggests that missing parts of a point cloud are likely to resemble the geometric structure of some objects in the same category. To exploit this similarity, we employ an adversarial learning strategy to learn the distinctive shape characteristics from category-specific data. Specifically, since depth maps provide a more compact and regular shape representation than partial point clouds and are easier to acquire, we render the reconstructed complete point cloud into depth maps from multiple viewpoints to capture the view-specific structures of the shape. Then, we utilize a 2D CNN-based network to discriminate the rendered depth maps and those from the same category in the dataset, thus refining our predictions based on category-specific geometric knowledge. Additionally, to address the anisotropic rendering issue, we propose a density-aware radius estimation algorithm to adaptively adjust the point radius from different viewpoints, thereby improving the quality of rendered images.
@@
Our MAL-SPC is unique in that it does not require multiple views of a partial point cloud for supervision or as additional data to drive the training process. Instead, it can achieve remarkable performance by employing only a single partial point cloud per object for training and supervision. Additionally, MAL-SPC effectively leverages the intrinsic geometry similarity from partial data to reconstruct missing structures. Moreover, unlike other self-supervised methods, our approach does not require any data preprocessing processes, such as random region removal, which is time-consuming and computationally intensive.
@@
We conduct extensive experiments on both synthetic datasets and real-world scans to demonstrate that our model significantly outperforms the state-of-the-art approaches by a large margin.
@@
  • We propose a Pattern Retrieval Network that leverages the intrinsic similarity and repetitiveness of object-level local geometry to reconstruct complete shapes. This network effectively infers the missing structures by learning correlated geometric patterns, including position and curvature, from the partial shape.
@@
The remainder of this paper is organized as follows. Section II briefly reviews existing supervised and unsupervised techniques for point cloud completion and reconstruction. Section III presents a detailed description of our proposed methods. Section IV discusses our experimental design, findings, and analysis. Finally, Section V concludes this paper.
@@ II-A Supervised Point Cloud Completion
The voxel-based methods [34, 35, 36, 37, 38, 39] first convert partial point clouds into 3D grids through voxelization and subsequently utilize the 3D Convolutional Neural Networks for learning the complete shapes. Dai et al. [34] employed a 3D-Encoder Predictor Network to generate an initial approximation of a shape, followed by a 3D shape synthesis method that refines the coarse output using a shape prior derived from a shape database. Xie et al. [35] introduced a Gridding module along with a reverse module for seamless conversion between point clouds and voxels, thereby minimizing structural information loss during the voxelization process. Additionally, they devised a Cubic Feature Sampling module to extract contextual features from neighboring points. Sun et al. [36] represented point clouds as occupancy grids and leveraged local patches of partial inputs to infer missing parts by capitalizing on the potential repetition of local patterns within an object. Initially, they employed a network to capture similar local patches of the inputs. Then, these patches undergo deformation and blending processes, ultimately reconstructing into the final outputs. Voxel-based methods can utilize the powerful feature extraction capabilities of CNNs. However, the computation speed of 3D CNN is relatively slow, and the conversion process between point clouds and voxels inevitably introduces accuracy loss.
@@ II-B Unsupervised Point Cloud Completion
Unsupervised point cloud methods do not require paired complete 3D ground truth to drive the completion process. These methods can be divided into three categories according to the supervision signal, i.e., unpaired methods, 2D-supervised methods, and self-supervised methods. The unpaired point cloud completion methods generate complete point clouds without paired ground truth as supervision during the training process [25, 26, 27, 28, 29, 30, 31, 29, 61]. In fact, these methods leverage complete and clean point clouds from other domains to train a model to capture 3D shape priors. Then, they used the learned shape priors to guide the completion of the partial point cloud from the source domain. The first unpaired method is PCL2PCL, proposed by Chen et al. [25]. PCL2PCL first used partial point clouds from the source domain to pre-train a partial-shape reconstruction Auto-encoder and then trained a complete-shape reconstruction Auto-encoder by utilizing complete point clouds from another domain. Subsequently, they employed a GAN-based model [62] to transform the partial global features to the same distribution with the complete shapes’ global features. To enhance the learning ability of the generation network, Wen et al. [26] proposed a cycle-based model to learn the transformation between partial and complete point clouds. Cao et al. [29] proposed a knowledge transfer model to learn the shape information of partial inputs. They designed a teacher-assistant-student network to transfer the knowledge learned from complete shapes to partial shapes. In addition to the shape priors provided by complete point clouds, Ma et al. [30] proposed to leverage the symmetry of objects to reconstruct the missing regions.
@@
The self-supervised methods only leverage the partial point cloud as a supervision signal to guide the completion process. Mittal et al. [63] proposed the first learning-based self-supervised method, PointPnCNet. PointPnCNet first randomly removed a small region from the input partial point cloud and used a network to predict the missing regions, including the manually removed parts and naturally missing parts. It employed the partial matching loss to preserve the existing parts in the predicted shapes. Inspired by closed-loop systems in control theory, Hong et al. [64] first predicted a coarse point cloud and used it to synthesize partial point clouds to input to the same model as the initial input. By utilizing consistency shape loss among predictions generated by multiple partial inputs, the model could generate complete shapes. Cui et al. [65] divided partial point clouds into three groups and then proposed a masked autoencoder to reconstruct complete shapes. Rather than directly using partial Chamfer Distance loss, they designed a Region-aware Chamfer Distance loss to reduce the noises of the predictions. Kim et al. [66] proposed a pose-aware approach for self-supervised point cloud completion. They projected the shape features onto various viewpoints and reconstructed the corresponding incomplete point clouds. They leveraged the shape consistency among multiple incomplete point clouds of the same object to constrain the reconstruction process.
@@ III Proposed Method
Let $\mathbf{D}_{0}\in\mathbb{R}^{H\times W}$ be the depth map scanned from an object. After back-projection, $\mathbf{D}_{0}$ yields a partial point cloud, denoted as $\mathbf{P}_{in}\in\mathbb{R}^{N\times 3}$. $\mathbf{P}_{com}\in\mathbb{R}^{M\times 3}$ represents the complete shape of $\mathbf{P}_{in}$, where $H$ and $W$ denote the height and width of the depth image, $N$ and $M$ are the numbers of points. Our goal is to find a mapping $f(\mathbf{P}_{in})=\mathbf{P}_{out}$ to make the predicted complete shape $\mathbf{P}_{out}$ as close as possible to $\mathbf{P}_{com}$ without $\mathbf{P}_{com}$ as supervision.
@@ III-A Pattern Retrieval Network
Since local geometric patterns of most objects exhibit similarities and repetitions, the missing regions of an object are likely to share similar or even identical local geometric patterns to the existing parts. To leverage this object-level geometric similarity, we propose a cross-attention-based pattern retrieval network to infer the missing structures.
As shown in Figure 2, consuming as input a partial point cloud $\mathbf{P}_{in}$, we first use a 3D encoder derived from PCN [15] to learn a global feature $\mathbf{o}$ that encapsulates the object’s shape information. Then we employ three layers of fully-connected layers to reconstruct a coarse yet complete point cloud $\mathbf{P}_{c}\in\mathbb{R}^{N\times 3}$. Next, we extract the position encodings and curvature encodings from the partial input and coarse output, respectively. The position encodings represent the positional relationship of local points, which are formulated as the average distance between each point and its nearest neighbors. Specifically, for each point $\mathbf{p}_{i}\in\mathbf{P}_{in}$, we first find out its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})=\{\mathbf{q}_{i}^{j}\}_{j=1}^{K}\subset\mathbf{P}_{in}$ and then compute the average Euclidean distance between $\mathbf{p}_{i}$ and $\mathcal{N}_{K}(\mathbf{p}_{i})$, i.e.,
@@
Figure 2: The flowchart of our Pattern Retrieval Network. “FC” and “MLP” denote the Fully-Connected layer and the Multi-Layer Perceptron. Given a partial point cloud $\mathbf{P}_{in}$, we first utilize an Auto-encoder to generate a coarse shape $\mathbf{P}_{c}$. Next, we separately extract position encodings and curvature encodings for both $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$, and then leverage a cross-attention-based module to identify and retrieve the most related encodings for each point in $\mathbf{P}_{c}$. Subsequently, $\mathbf{P}_{c}$ and the selected encodings are upsampled and fused to refine the shape.
@@
Figure 3: Depth maps rendered from point clouds with an equal point number. The rendering process employs a fixed and identical point radius. It is obvious that, with a small foreground area, the visible surface can be well represented. However, when the foreground area is large, a fixed point radius cannot cover the entire visible surface, resulting in some foreground pixels being covered by invisible points.

### v2

@@ Abstract
In real-world scenarios, scanned point clouds are often incomplete due to occlusion issues. The task of self-supervised point cloud completion involves reconstructing missing regions of these incomplete objects without the supervision of complete ground truth. Current self-supervised methods either rely on multiple views of partial observations for supervision or overlook the intrinsic geometric similarity that can be identified and utilized from the given partial point clouds. In this paper, we propose MAL-SPC, a framework that effectively leverages both region-level and category-specific geometric similarities to complete missing structures. Our MAL-SPC does not require any 3D complete supervision and only necessitates a single partial point cloud for each object. Specifically, we first introduce a Pattern Retrieval Network to retrieve similar position and curvature patterns between the partial input and the predicted shape, then leverage these similarities to densify and refine the reconstructed results. Additionally, we render the reconstructed complete shape into multi-view depth maps and design an adversarial learning module to learn the geometry of the target shape from category-specific single-view depth images. To achieve anisotropic rendering, we design a density-aware radius estimation algorithm to improve the quality of the rendered images. Our MAL-SPC yields the best results compared to current state-of-the-art methods. We will make the source code publicly available at https://github.com/ltwu6/malspc.
@@ I Introduction
Point clouds, which represent instances as a collection of points in 3D space, have attracted increasing interest due to the growing popularity of 3D scanning devices such as LiDAR and depth cameras. They offer a detailed representation of the geometric features and spatial structures of 3D objects and scenes, and are widely applied in various fields, such as autonomous driving[1, 2, 3, 4, 5], motion analysis[6, 7], and augmented reality[8, 9]. Accurate and complete shape representation plays a vital role in understanding 3D point clouds. However, point clouds scanned in practical scenarios are often incomplete due to occlusion. This incompleteness severely impacts the performance of downstream tasks, such as 3D object detection [10, 11, 12] and scene understanding [13, 14].
@@
To solve this problem, point cloud completion methods have emerged. Point cloud completion aims to infer the missing regions of an object based on the partial point cloud. Most of the existing point cloud completion methods rely on a large number of complete point clouds as supervision [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. However, in real-world scenes, clean and complete point clouds are really few since current devices struggle to directly capture such data.
@@
To overcome these limitations, some research has focused on unsupervised learning approaches for point cloud completion. Among these methods, some employ complete shapes unpaired with the partial input as priors to drive the completion process [25, 26, 27, 28, 29, 30, 31], while others utilize 2D signals such as RGB images or silhouette maps from multiple viewpoints as supervision [32, 33]. However, the former still requires complete point clouds to acquire shape priors, and the latter relies on multi-view images along with the projection parameters. Recently, self-supervised methods, which only require partial inputs for supervision, have been proposed. These methods estimate the missing parts of the point cloud by utilizing the multi-view consistency of partial scans and the robust learning capabilities of the masked models. However, they face some limitations. First, these methods rely on multi-view partial point clouds of each object as shape constraints or as training data to drive the shape learning process. Nevertheless, obtaining point clouds from multiple viewpoints is also impractical in real-world applications. Secondly, they fail to utilize the inherent local similarities and category-specific geometric properties of objects.
@@
From our perspective, it is commonly observed that an object frequently exhibits similar or repetitive local structures. For instance, a table typically features four identical cylindrical legs. This observation implies that the existing and missing sections of a point cloud are likely to share similar local geometric patterns. Moreover, objects within the same category generally display consistent structures. For example, most airplanes feature two flat wings. Consequently, these inherent geometric similarities, specific to each category, can be utilized to guide the completion of missing regions.
@@
In this paper, we propose MAL-SPC, a Multi-view Adversarial Learning framework for Self-supervised Point cloud Completion. One of the core components, the Pattern Retrieval Network (PRN), reconstructs complete shapes by initially generating a coarse shape and subsequently encoding local geometric patterns from both the coarse and partial input shapes. These patterns manifest in two forms: position encodings, quantified as the average distance between each point and its nearest neighbors, and curvature encodings, defined by the variations in normals between each point and its closest neighbors. Subsequently, a cross-attention-based module is introduced to retrieve correlated local patterns for the coarse shape from the partial shape. These retrieved patterns are further utilized to guide the densification and refinement of the coarse shape, capitalizing on the region-level geometric similarities and repetitiveness inherent in the data.
@@
In addition, we introduce a Multi-view Adversarial Network (MAN) to learn the shared geometry within each category. Objects within the same category commonly exhibit remarkably similar shape distributions because they share many typical geometric structures. This similarity suggests that missing regions of a point cloud are likely to resemble the geometric structure of some objects in the same category. To exploit this similarity, we employ an adversarial learning strategy to learn the distinctive shape characteristics from category-specific data. Specifically, since depth maps provide a more compact and regular shape representation and are easier to acquire than partial point clouds, we render the reconstructed complete point cloud into depth maps from multiple viewpoints to capture the view-specific structures of the shape. Then, we utilize a 2D CNN-based network to discriminate the rendered depth maps and those from the same category in the dataset, thus refining our predictions based on category-specific geometric knowledge. Additionally, to address the anisotropic rendering issue, we propose a density-aware radius estimation algorithm to adaptively adjust the point radius from different viewpoints, thereby improving the quality of rendered images.
@@
Our MAL-SPC is unique in that it does not require multiple views of a partial point cloud for supervision or as additional data to drive the training process. Instead, it can achieve remarkable performance by employing only a single partial point cloud per object for training and supervision. Additionally, MAL-SPC effectively leverages the intrinsic geometry similarity from partial data to reconstruct missing structures. Furthermore, unlike other self-supervised methods, our approach does not require any data preprocessing processes, such as random region removal, which is time-consuming and computationally intensive.
@@
We conduct extensive experiments on both synthetic datasets and real-world scans. The experimental results demonstrate that our model outperforms the state-of-the-art approaches.
@@
  • We propose a Pattern Retrieval Network that leverages the intrinsic similarity and repetitiveness of region-level local geometry to reconstruct complete shapes. This network effectively infers the missing structures by learning correlated geometric patterns, including position and curvature, from the partial shape.
@@
The remainder of this paper is organized as follows. Section II briefly reviews existing supervised and unsupervised techniques for point cloud completion and reconstruction. Section III provides a detailed description of our proposed methods. Section IV discusses our experimental design, findings, and analysis. Section V presents the limitations of our approach and future work. Finally, Section VI concludes this paper.
@@ II-A Supervised Point Cloud Completion
The voxel-based methods [34, 35, 36, 37, 38, 39, 40] first convert partial point clouds into 3D grids through voxelization and subsequently utilize the 3D Convolutional Neural Networks for learning the complete shapes. Dai et al. [34] employed a 3D-Encoder Predictor Network to generate an initial approximation of a shape, followed by a 3D shape synthesis method that refines the coarse output using a shape prior derived from a shape database. Xie et al. [35] introduced a Gridding module along with a reverse module for seamless conversion between point clouds and voxels, thereby minimizing structural information loss during the voxelization process. Additionally, they devised a Cubic Feature Sampling module to extract contextual features from neighboring points. Sun et al. [36] represented point clouds as occupancy grids and leveraged local patches of partial inputs to infer missing parts by capitalizing on the potential repetition of local patterns within an object. Initially, they employed a network to capture similar local patches of the inputs. Then, these patches undergo deformation and blending processes, ultimately reconstructing into the final outputs. Voxel-based methods can utilize the powerful feature extraction capabilities of CNNs. However, the computation speed of 3D CNN is relatively slow, and the conversion process between point clouds and voxels inevitably introduces accuracy loss.
@@ II-B Unsupervised Point Cloud Completion
Unsupervised point cloud methods do not require paired complete 3D ground truth to drive the completion process. These methods can be divided into three categories according to the supervision signal, i.e., unpaired methods, 2D-supervised methods, and self-supervised methods. The unpaired point cloud completion methods generate complete point clouds without paired ground truth as supervision during the training process [25, 26, 27, 28, 29, 30, 31, 29, 62, 63]. In fact, these methods leverage complete and clean point clouds from other domains to train a model to capture 3D shape priors. Then, they used the learned shape priors to guide the completion of the partial point cloud from the source domain. The first unpaired method is PCL2PCL, proposed by Chen et al. [25]. PCL2PCL first used partial point clouds from the source domain to pre-train a partial-shape reconstruction Auto-encoder and then trained a complete-shape reconstruction Auto-encoder by utilizing complete point clouds from another domain. Subsequently, they employed a GAN-based model [64] to transform the partial global features to the same distribution with the complete shapes’ global features. To enhance the learning ability of the generation network, Wen et al. [26] proposed a cycle-based model to learn the transformation between partial and complete point clouds. Cui et al. [63] proposed a latent-space energy-based model to learn a conditional distribution of the complete shape encoding given the partial encoding. Then, they designed a residual sampling strategy to select a valid encoding and decode it into a complete shape. Cao et al. [29] proposed a knowledge transfer model to learn the shape information of partial inputs. They designed a teacher-assistant-student network to transfer the knowledge learned from complete shapes to partial shapes. In addition to the shape priors provided by complete point clouds, Ma et al. [30] proposed to leverage the symmetry of objects to reconstruct the missing regions.
@@
The self-supervised methods only leverage the partial point cloud as a supervision signal to guide the completion process. Mittal et al. [65] proposed the first learning-based self-supervised method, PointPnCNet. PointPnCNet first randomly removed a small region from the input partial point cloud and used a network to predict the missing regions, including the manually removed parts and naturally missing parts. It employed the partial matching loss to preserve the existing parts in the predicted shapes. However, PointPnCNet relies on multiple partial observations for supervision, whereas our approach requires only a single-view partial point cloud. Inspired by closed-loop systems in control theory, Hong et al. [66] first predicted a coarse point cloud and used it to synthesize partial point clouds, and then input them to the same model as the initial input. By utilizing consistency shape loss among predictions generated by multiple partial inputs, the model could generate complete shapes. Nevertheless, this method requires the same training on the testing objects. Cui et al. [67] divided partial point clouds into three groups and then proposed a masked autoencoder to reconstruct complete shapes. Rather than directly using partial Chamfer Distance loss, they designed a Region-aware Chamfer Distance loss to reduce the noises of the predictions. Kim et al. [68] proposed a pose-aware approach for self-supervised point cloud completion. They projected the shape features onto various viewpoints and reconstructed the corresponding incomplete point clouds. They leveraged the shape consistency among multiple incomplete point clouds of the same object to constrain the reconstruction process.
@@ III Proposed Method
Let $\mathbf{D}_{0}\in\mathbb{R}^{H\times W}$ be the depth map scanned from an object. After back-projection, $\mathbf{D}_{0}$ yields a partial point cloud, denoted as $\mathbf{P}_{in}\in\mathbb{R}^{N\times 3}$. Denote by $\mathbf{P}_{com}\in\mathbb{R}^{M\times 3}$ the complete shape of $\mathbf{P}_{in}$, where $H$ and $W$ represent the height and width of the depth image, $N$ and $M$ are the numbers of points. Our goal is to find a mapping $f(\mathbf{P}_{in})=\mathbf{P}_{out}$ to make the predicted complete shape $\mathbf{P}_{out}$ as close as possible to $\mathbf{P}_{com}$ without $\mathbf{P}_{com}$ as supervision.
@@ III-A Pattern Retrieval Network
The Pattern Retrieval Network leverages objects’ region-level geometric similarity to infer the missing structures. As shown in Figure 2, consuming as input a partial point cloud $\mathbf{P}_{in}$, we first use a 3D encoder derived from PCN [15] to learn a global feature $\mathbf{o}$ that encapsulates the object’s shape information. Then we employ three layers of fully-connected layers to reconstruct a coarse yet complete point cloud $\mathbf{P}_{c}\in\mathbb{R}^{N\times 3}$. Next, we extract the position encodings and curvature encodings from the partial input and coarse output, respectively. The position encodings represent the positional relationship of local points, which are formulated as the average distance between each point and its nearest neighbors. Specifically, for each point $\mathbf{p}_{i}\in\mathbf{P}_{in}$, we first find out its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})=\{\mathbf{q}_{i}^{j}\}_{j=1}^{K}\subset\mathbf{P}_{in}$ and then compute the average Euclidean distance between $\mathbf{p}_{i}$ and $\mathcal{N}_{K}(\mathbf{p}_{i})$, i.e.,
@@
Figure 2: The flowchart of our Pattern Retrieval Network. “FC” and “MLP” denote the Fully-Connected layer and the Multi-Layer Perceptron layer. Given a partial point cloud $\mathbf{P}_{in}$, we first utilize an Auto-encoder to generate a coarse shape $\mathbf{P}_{c}$. Next, we separately extract position encodings and curvature encodings for both $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$, and then leverage a cross-attention-based module to identify and retrieve the most related encodings for each point in $\mathbf{P}_{c}$. Subsequently, $\mathbf{P}_{c}$ and the selected encodings are upsampled and fused to refine the shape.
@@
Figure 3: Depth maps rendered from point clouds with an equal point number. The rendering process employs a fixed and identical point radius. With a small foreground area, the visible surface can be well represented. However, when the foreground area is large, a fixed point radius cannot cover the entire visible surface, resulting in some foreground pixels being covered by invisible points.